# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v2

@@ VII.2 Dilution of asymmetry
i.e., $\mathcal{A}_{\mathrm{c}}(\rho)=1/R(\rho\to\phi)$, where $1/\infty$ and $1/0$ are formally regarded as 00 and $\infty$, respectively. We remark that Eq. (8) implies $\mathcal{A}_{\mathrm{c}}(\rho)<\infty$ only if $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$. Thus, we only investigate $\mathcal{A}_{\mathrm{c}}(\rho)$ for $\rho$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$.
@@
where $\{\rho_{i}\}_{i=1}^{k}$ is a set of states and $\{p_{i}\}_{i=1}^{k}$ is a probability distribution such that $p_{i}>0$ and $\sum_{i=1}^{k}p_{i}=1$. An intuition behind this inequality is that, instead of preparing $\rho$ directly, we can adopt a strategy where $\{\rho_{i}\}_{i=1}^{k}$ is probabilistically generated according to the probability distribution $\{p_{i}\}_{i=1}^{k}$, although this strategy is not necessarily optimal. See Appendix E for the proof of Eq. (75). We remark that the left-hand side of Eq. (75) is finite only if $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho_{i})$ for all $i=1,\cdots,k$.
In light of the pure-state conversion theory established in Theorem 1, a promising approach to obtaining a computable upper bound of $\mathcal{A}_{\mathrm{c}}(\rho)$ is to decompose a state $\rho$ into an ensemble of pure states. Suppose that a state $\rho$ can be decomposed as $\rho=\sum_{i=1}^{k-1}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$, where a symmetric (possibly mixed  [^2] We remark that not all symmetric mixed states can be decomposed as an ensemble of symmetric pure states. Indeed, when a unitary representation of a group is decomposed into irreducible components, Schur’s lemma implies that only a one-dimensional irreducible representation contains a symmetric pure state. In other words, no symmetric pure state exists in any higher-dimensional irreducible representation.) state $\rho_{\mathrm{s}}$ and pure states $\{\psi_{i}\}_{i=1}^{k-1}$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\psi_{i})$. Then, Eq. (7) implies
@@
When a state $\rho$ can be obtained as an ensemble of pure states and symmetric state, minimizing the left-hand side of Eq. (76) over such an ensemble leads to a tighter bound on the asymmetry cost $\mathcal{A}_{\mathrm{c}}(\rho)$. In the case of $G=U(1)$, it is known that such a minimization yields a tight bound [21]. Let $U(\theta)=e^{\mathrm{i}H\theta}$ and $U^{\prime}(\theta)=e^{\mathrm{i}H^{\prime}\theta}$ be unitary representations of $G=U(1)$ on finite-dimensional Hilbert spaces with Hamiltonians $H=\sum_{n=0}^{d-1}n\ket{n}\bra{n}$ | and $H^{\prime}=\sum_{n=0}^{d^{\prime}-1}n\ket{n}\bra{n}$ |. We fix the reference state $\phi$ to be the so-called coherence bit, given by |$\ket{\phi}=(\ket{0}+\ket{1})/\sqrt{2}$ and assume that $\mathrm{Sym}_{G}(\rho)=\mathrm{Sym}_{G}(\phi)$. Since $r_{i}=V(\psi_{i},H^{\prime})/V(\phi,H)=4V(\psi_{i},H^{\prime})$, Eq. (76) implies $\min_{\{p_{i},\psi_{i}\}_{i}}4\sum_{i}p_{i}V(\psi_{i},H^{\prime})\geq\mathcal{A}_{\mathrm{c}}(\rho)$, where the minimization is taken over the set of all pure-state ensembles $\{p_{i},\psi_{i}\}_{i}$ satisfying $\mathrm{Sym}_{G}(\rho)\subset\mathrm{Sym}_{G}(\psi_{i})$ and $\rho=\sum_{i}p_{i}\psi_{i}$. Note that it suffices to consider pure-state ensembles for $G=U(1)$ since any symmetric state is given as a probabilistic mixture of symmetric pure states. It was conjectured in [74] and later proven in [75] that $\min_{\{p_{i},\psi_{i}\}_{i}}4\sum_{i}p_{i}V(\psi_{i},H^{\prime})=\mathcal{F}_{H^{\prime}}(\rho)$, where $\mathcal{F}_{H^{\prime}}(\rho)$ denotes the SLD quantum Fisher information, given by $\mathcal{F}_{H^{\prime}}(\rho)\coloneqq 2\sum_{i,j}\frac{(\lambda_{i}-\lambda_{j})^{2}}{\lambda_{i}+\lambda_{j}}|\braket{i}{H}{j}|^{2}$ using the eigenvalue decomposition $\rho=\sum_{i}\lambda_{i}\ket{i}\bra{i}$ |. Furthermore, in [21], it was shown that the minimum is attained by a set of pure states $\{\psi_{i}\}_{i}$ such that $\mathrm{Sym}_{G}(\rho)\subset\mathrm{Sym}_{G}(\psi_{i})$, which implies $\mathcal{F}_{H^{\prime}}(\rho)\geq\mathcal{A}_{\mathrm{c}}(\rho)$. In [21], the opposite inequality was also shown by proving the monotonicity of the SLD quantum Fisher information rate under asymptotic conversion, which completed the proof of $\mathcal{F}_{H^{\prime}}(\rho)=\mathcal{A}_{\mathrm{c}}(\rho)$ for $G=U(1)$. In contrast to the $U(1)$ case, however, we remark that decomposition into $\rho=\sum_{i=1}^{k-1}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$ with symmetric $\rho_{\mathrm{s}}$ and pure states $\{\psi_{i}\}_{i}$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\psi_{i})$ does not always exist for a general Lie group $G$. See Appendix E for a concrete example for $G=U(4)$.

### v3

@@ VII.2 Dilution of asymmetry
i.e., $\mathcal{A}_{\mathrm{c}}(\rho)=1/R(\rho\to\phi)$, where $1/\infty$ and $1/0$ are formally regarded as 0 and $\infty$, respectively. We remark that Eq. (9) implies $\mathcal{A}_{\mathrm{c}}(\rho)<\infty$ only if $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$. Thus, we only investigate $\mathcal{A}_{\mathrm{c}}(\rho)$ for $\rho$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$.
@@
where $\{\rho_{i}\}_{i=1}^{k}$ is a set of states and $\{p_{i}\}_{i=1}^{k}$ is a probability distribution such that $p_{i}>0$, $\sum_{i=1}^{k}p_{i}=1$, and $\sum_{i=1}^{k}p_{i}\rho_{i}=\rho$. An intuition behind this inequality is that, instead of preparing $\rho$ directly, we can adopt a strategy where $\{\rho_{i}\}_{i=1}^{k}$ is probabilistically generated according to the probability distribution $\{p_{i}\}_{i=1}^{k}$, although this strategy is not necessarily optimal. See Appendix E for the proof of Eq. (79). We remark that the left-hand side of Eq. (79) is finite only if $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho_{i})$ for all $i=1,\cdots,k$.
In light of the pure-state conversion theory established in Theorem 1, a promising approach to obtaining a computable upper bound of $\mathcal{A}_{\mathrm{c}}(\rho)$ is to decompose a state $\rho$ into an ensemble of pure states. Suppose that a state $\rho$ can be decomposed as $\rho=\sum_{i=1}^{k-1}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$, where $\rho_{\mathrm{s}}$ is a symmetric (possibly mixed  [^2] We remark that not all symmetric mixed states can be decomposed as an ensemble of symmetric pure states. Indeed, when a unitary representation of a group is decomposed into irreducible components, Schur’s lemma implies that only a one-dimensional irreducible representation contains a symmetric pure state. In other words, no symmetric pure state exists in any higher-dimensional irreducible representation.) state, and $\{\psi_{i}\}_{i=1}^{k-1}$ are pure states satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\psi_{i})$. Then, Eq. (8) implies
@@
Similarly to Eq. (11), this quantity can be expressed by using the quantum max-relative entropy as
$r_{i}=2^{\min_{j}D_{\max}\left(\mathcal{Q}^{\mathcal{U}_{g_{j}}^{\prime}(\psi_{i})}\middle\|\mathcal{Q}^{\mathcal{U}_{g_{j}}(\phi)}\right)},$ (82)
where $g_{j}\in G_{j}$ is a representative element of each connected component $G_{j}$ of $G$.
[…]
In contrast to the $U(1)$ case, a decomposition into $\rho=\sum_{i=1}^{k-1}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$ with symmetric $\rho_{\mathrm{s}}$ and pure states $\{\psi_{i}\}_{i}$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\psi_{i})$ does not always exist for a general Lie group $G$. See Appendix E for a concrete example for $G=U(4)$. In such cases, Eq. (80) does not provide a meaningful upper bound on the asymmetry cost. However, by carefully refining the estimate-and-convert approach introduced in Sec. V.3, we find that this limitation due to the symmetry subgroups can be circumvented. Specifically, in Appendix F, we show the following:
Proposition 14 (Upper bound on the asymmetry cost).
For a reference state $\phi$, consider a state $\rho$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$. The asymmetry cost of a state $\rho$ is upper bounded as follows:
$\min\sum_{i}p_{i}r_{i}\geq\mathcal{A}_{\mathrm{c}}(\rho),$ (83)
where $r_{i}$ is defined in Eq. (81), and the minimization is taken over all the decompositions of $\rho$ into ensemble $\rho=\sum_{i}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$, such that $\{\psi_{i}\}_{i}$ are pure states, $\rho_{\mathrm{s}}$ is a symmetric state, and $\{p_{i}\}_{i},p_{\mathrm{s}}$ denote a probability distribution satisfying $p_{i},p_{\mathrm{s}}\geq 0$ and $\sum_{i}p_{i}+p_{\mathrm{s}}=1$.